# Paper2Code: Automating Code Generation from Scientific Papers in Machine Learning

### v2

@@ 3.2.3 Coding
The final stage of our methodology is the coding phase, which produces code constituting the research repository. The generation of each file is guided by the comprehensive outputs from previous phases: the research paper itself, the overall plan, architecture design, logic designs, the configuration file, file-specific analyses, and previously generated code. As repository files often exhibit import dependencies among themselves, we adhere strictly to the ordered file list established during the planning phase, ensuring sequential consistency. Formally, the file-specific code generation process is defined as $\{M_{\text{coder}}(R,P=\{o,d,l,g\},f_{i},a_{i})=c_{i}\}_{i=1}^{n=|\text{file}|};\text{where }f\in F,a\in A,c\in C$. This process is applied sequentially for each file to ensure a structured and logically coherent repository implementation $c_{i}$. The initially produced code may require subsequent debugging or refinement to ensure correctness and full functionality. In this work, comprehensive debugging strategies and detailed error-correction workflows remain beyond the current scope of this paper.
@@ 4 Experimental Setup
4.1 Data
Paper2Code Benchmark
To evaluate the effectiveness of our framework for reproducing both the proposed methods and experiments described in scientific papers, we constructed a experimental benchmark based on papers from ICML 2024, NeurIPS 2024, and ICLR 2024. Using the OpenReview API [^1] https://docs.openreview.net/reference/api-v2, we first filtered the accepted papers from each conference to include only those with publicly available GitHub repositories. Among these, we selected repositories with a total codebase of fewer than 70,000 tokens to ensure reproducibility within a manageable scope. To ensure quality, we performed model-based evaluations of the repositories and selected the top 30 highest-scoring papers from each venue. These 90 papers constitute our Paper2Code benchmark for experimental validation. A detailed list of the 90 papers is provided in Table 9, Table 10 and Table 11. For the human evaluation, we select 13 papers authored by the evaluation participants, and listed in Table 13.
PaperBench Code-Dev
We further validate our framework using the PaperBench Code-Dev benchmark [40], which consists of a curated set of 20 papers from ICML 2024.
@@ 4.2 Baselines and Our Model
Paper2Code Benchmark
To the best of our knowledge, there is no existing framework that directly addresses the task of end-to-end paper-to-code reproduction. However, we consider several related approaches as baselines, particularly those designed for software development from natural language inputs such as requirements. These serve as the closest comparison points in terms of producing functional code repositories. In addition, we include comparisons with a number of ablated variants that serve as simple or naive baselines, as follow: 1. ChatDev [36] - It is a multi-agent framework where LLM-driven agents collaboratively develop software through dialogues. In our experiments, the entire paper is provided as input(requirements) to generate a full code repository; 2. MetaGPT [15] - It adopts a role-based multi-agent paradigm, where software development is organized via Standardized Operating Procedures (SOPs), which specify structured workflows and responsibilities for each agent. We provide the entire paper as input and prompt the system to construct the full code repository; 3. Abstract - A naive baseline where only the paper’s abstract is provided to the language model, requiring it to implement the code repository based on minimal information; 4. Paper - Another naive baseline in which the full paper is provided as input, and the model is prompted to generate a corresponding code repository; 5. PaperCoder (Ours) - Our proposed method, which integrates planning, analysis and code generation to generate high-quality repositories from scientific papers; 6. Oracle - The official code repository released by the paper’s authors. This serves as an upper bound for comparison, reflecting the intended implementation of the proposed in the paper.
PaperBench Code-Dev
We compare our method against the two agent variants reported in the original benchmark. The Basic Agent follows a ReAct-style approach and is equipped with a predefined set of tools, including a bash shell command executor, a Python code runner, a web browser, and a paginated file reader for processing long documents. The Iterative Agent extends the Basic Agent by utilizing the full available computation time and incorporating custom prompting strategies that encourage step-by-step reasoning through subgoal decomposition (see Starace et al. [40] for details).
4.3 Evaluation Setups
4.3.1 Paper2Code Benchmark
Evaluating code generation typically relies on unit tests or verifying specific implementation components [5, 11]. However, such evaluations are often infeasible due to the lack of gold-standard repositories. Many papers do not release official code, and even when they do, test scripts are frequently unavailable. Additionally, manually annotating each paper with ground-truth implementation details is highly labor-intensive, rendering large-scale evaluation impractical.
Model-Based Evaluation
To address this challenge, we adopt a model-based evaluation approach [50, 13, 24] that assesses repository quality with or without access to gold-standard repository. We consider two variants: (1) reference-based, which leverages both the paper and the author-provided repository, and (2) reference-free, which evaluates the generated repository using only the paper. For both variants, we prompt the language model to critique the required implementation components and evaluate their correctness. It assigns severity levels (high, medium and low) to any missing or flawed elements and generates an overall correctness score on a 1 to 5 scale. To ensure stability, we report the average score across multiple generations using $n$-way sampling. In this experiments, we set the $n=8$ and use the o3-mini-high as a evaluation model.
@@ 4.3.1 Paper2Code Benchmark
When an official repository is available, we treat it as one plausible implementation of the method described in the paper. While multiple valid implementations may exist, the author-released version is generally considered the most accurate, as it best captures the core components necessary for reproducing experimental results. In this setting, we provide both the paper and the gold-standard repository as input, and prompt the evaluation model to identify and critique the required components. The model then compares the predicted repository against these components and assigns a correctness score from 1 to 5, reflecting both component coverage and the severity of any errors.
@@
In many cases, author-released code is unavailable. To handle such scenarios, we propose a reference-free evaluation strategy that relies solely on the paper and the generated repository. As in the reference-based setting, the model is prompted to infer the required implementation components from the paper, critique them, and assess whether they are adequately implemented in the generated repository. A correctness score from 1 to 5 is then assigned based on the presence and quality of these components.
@@
[…]
4.3.2 PaperBench Code-Dev Evaluation
For evaluating our system on the PaperBench Code-Dev benchmark, we adopt the official evaluation protocol, which measures replication accuracy across a curated set of ICML 2024 papers. Specifically, we follow the rubric authored by the original paper’s authors, which defines a hierarchical set of implementation requirements. The evaluation model assesses each generated repository based on whether the submitted code correctly implements any of the specified requirements. Importantly, the evaluation focuses only on the code development node which whether the candidate repository contains a correct implementation of some requirement.
Table 1: Results on our experimental Paper2Code benchmark. We report average model-based evaluation scores for each conference. Oracle denotes the official repository released by the paper’s authors. Reference-based evaluation assesses correctness (on a 1–5 scale) by comparing the generated repository against both the paper and the official implementation, while reference-free evaluation relies solely on the paper. We report statistics on the average number of tokens, files, and functions per repository. The best scores are highlighted in bold.
Reference-based Reference-free Statistics
ICML NeurIPS ICLR ICML NeurIPS ICLR # of Tokens # of Files # of Funcs
ChatDEV 2.97 (0.58) 2.96 (0.69) 2.70(0.63) 4.12 (0.53) 4.01 (0.74) 4.00 (0.65) 6150.54 6.99 23.82
MetaGPT 2.75 (0.70) 2.95 (0.87) 2.48 (0.48) 3.63 (0.75) 3.59 (0.92) 3.52 (0.60) 5405.21 3.24 18.08
Abstract 2.43 (0.49) 2.35 (0.62) 2.28 (0.42) 3.01 (0.60) 2.99 (0.78) 3.03 (0.64) 3376.99 1.28 12.62
Paper 3.28 (0.67) 3.22 (0.80) 3.08 (0.66) 4.30 (0.53) 4.08 (0.84) 4.15 (0.63) 3846.33 1.79 14.84
PaperCoder 3.72 (0.54) 3.83 (0.50) 3.68 (0.52) 4.73 (0.44) 4.77 (0.38) 4.73 (0.32) 14343.38 6.97 35.22
Oracle - - - 4.80 (0.32) 4.83 (0.38) 4.84 (0.26) 32149.04 28.00 122.03
@@ 5 Experimental Results and Analysis
Table 1 presents the main experimental results, our method consistently outperform across all conferences and both evaluation modes. Under the reference-based setting, our approach achieves the highest average correctness scores of 3.72, 3.83, and 3.68 on ICML, NeurIPS, and ICLR papers, respectively. Similarly, in the reference-free evaluation, our method attains scores of 4.73, 4.77, and 4.73, outperforming all baselines. Compared to software development work baselines such as ChatDev and MetaGPT, our method exhibits substantial performance gains. Notably, although ChatDev generates a comparable number of files (6.99 vs. 6.97), our method produces significantly more functions (35.22 vs. 23.82), indicating a higher level of granularity and completeness in the generated repositories. In contrast, MetaGPT lags behind both in evaluation scores and in code quantity metrics. Ablation variants using only the abstract or full paper underperform relative to ours. This highlights the effectiveness of our multi-stage framework in systematically extracting and organizing implementation-relevant information. For reference, we include an Oracle score that represents the upper bound, corresponding to the official implementation by the authors. While there is still a gap to the Oracle, our method achieves comparable results that are closer to this upper bound compared to existing LLM-based frameworks. Overall, these results demonstrate that our proposed system not only surpasses prior approaches in generating accurate and comprehensive code repositories but also excels in producing fine-grained, detail-oriented implementations necessary for faithful reproduction of research methods.
@@
Figure 3: Correlation between reference-based and reference-free model-based evaluations.
Correlation between Reference-based and Reference-free
We investigate the alignment between reference-based and reference-free evaluation metrics by computing their correlation across a large set of generated repositories. As shown in Figure 3, the two evaluation scores exhibit a strong positive correlation, with a Pearson correlation coefficient of $r=0.79$ and a statistically significant $p$-value of $0.00$. This high correlation suggests that reference-free evaluation which does not rely on access to ground-truth code can serve as a reliable proxy for reference-based assessment. The result supports the feasibility of using reference-free methods in scenarios where official implementations are unavailable, while still maintaining consistency with traditional evaluation standards.
Table 2: Results on the PaperBench Code-Dev benchmark using o3-mini-high. For BasicAgent and IterativeAgent, results are averaged over three runs, and standard errors are reported.
Model Replication scores (%)
BasicAgent $5.1\pm 0.8$
IterativeAgent $16.4\pm 1.4$
PaperCoder 44.26
PaperBench Code-Dev Results
To assess the effectiveness of our approach in realistic research scenarios, we evaluate it on the PaperBench Code-Dev benchmark. As shown in Table 2, our method achieves a best replication score. This demonstrates the superiority of our structured multi-agent pipeline, which performs coordinated planning, analysis, and implementation.
Table 3: Human evaluation results. For model-based evaluations, including reference-based and reference-free settings, correctness scores are converted into rankings for comparability. In the human evaluation, 13 authors were asked to rank repositories within two groups based on method type. Rankings are then converted into scores by assigning 5, 3, and 1 points to the first-, second-, and third-ranked repositories, respectively.
@@
Abstract 2.36 (0.34) 2.99 (0.53) 1.62 (1.26) 3.00 (0.00) 3.00 (0.00) 2.69 (0.63)
Paper 3.19 (0.44) 4.22(0.52) 3.15 (1.28) 1.86 (0.36) 1.79 (0.43) 1.92 (0.64)
PaperCoder (Ours) 3.74 (0.30) 4.71 (0.27) 4.23 (1.30) 1.14 (0.36) 1.07 (0.27) 1.38 (0.65)
ChatDev 2.67(0.63) 3.87 (0.36) 2.69 (1.11) 2.50 (0.52) 2.36 (0.50) 2.15 (0.55)
MetaGPT 2.65 (0.46) 3.38 (0.66) 1.77 (1.30) 2.07 (0.52) 2.14 (0.52) 2.61 (0.65)
PaperCoder (Ours) 3.74 (0.30) 4.71 (0.27) 4.54 (1.20) 1.00 (0.00) 1.00 (0.00) 1.23 (0.60)
@@
To assess the practical utility and subjective quality of the generated code, we conduct a comprehensive human evaluation. We include both model-based evaluations (reference-based and reference-free) and direct human assessments, where authors or expert annotators rank the quality of each generated repository. Table 3 summarizes the results across these three evaluation types. Across all evaluation criteria, our method consistently achieves the highest scores and the best rankings. In terms of average human evaluation scores, our system achieves 4.23 and 4.38 comparing both ablation variations and software development works, respectively, substantially outperforming all baselines. Similarly, in model-based reference-based and reference-free evaluations, our method attains 3.74 (ref-based) and 4.71 (ref-free), outperforming alternatives such as ChatDEV and MetaGPT. In the ranking-based evaluation, our method also secures the top position across all metrics. Notably, it achieves the best average rank in human evaluations (1.38 and 1.31), while other baselines such as MetaGPT and only Abstract consistently rank lower. The low variance in our ranking (e.g., 1.00 in both reference-based and reference-free) further demonstrates the robustness of our system across diverse evaluation settings. These results clearly indicate that our proposed approach produces more useful and accurate implementations than existing methods, regardless of whether the evaluation is automated or conducted by human experts.
Table 4: Human evaluation results on model variants of PaperCoder. DS-Coder refers to DeepSeek-Coder-V2-Lite-Instruct, Qwen-Coder to Qwen2.5-Coder-7B-Instruct, and DS-Distil-Qwen to DeepSeek-R1-Distill-Qwen-14B. o3-mini-high denotes the high reasoning-effort of o3-mini. Other settings follow Table  3.
DS-Coder Qwen-Coder DS-Distill-Qwen o3-mini-high
Score ($\uparrow$) Ref-based 1.63 (0.43) 1.8 (0.28) 2.07 (0.30) 3.74 (0.30)
Ref-free 1.82 (0.39) 2.1 (0.28) 2.29 (0.29) 4.71 (0.27)
Human 1.41 (0.64) 2.74 (1.14) 3.05 (1.04) 4.79 (0.74)
Ranking ($\downarrow$) Ref-based 3.36 (0.93) 2.93 (0.62) 2.36 (0.63) 1.00 (0.00)
Ref-free 3.36 (0.84) 2.86 (0.66) 2.14 (0.36) 1.00 (0.00)
Human 3.69 (0.48) 2.69 (0.85) 2.46 (0.78) 1.15 (0.55)
Detailed Analysis on Generated Repository
We conduct a fine-grained human evaluation to better understand the practical utility and implementation quality of the generated repositories. Specifically, we asked authors to review a set of repositories about covering both baselines and ablated variants and to select the one they considered most suitable for reproducing their work. 77% (10 out of 13 authors) selected the repository generated by our method as the top choice, followed by 3 authors who preferred the Paper variant. The most common reasons cited for favoring our method include completeness, clean structure, and faithfulness to the original paper. Full results are shown in Table 8.
Figure 4: Human evaluation results on the helpfulness and section-level implementation accuracy of repositories generated by our framework.
[…]
Table 5: Rank correlation between model-based and human evaluations is measured using GPT-4o (gpt-4o-2024-11-20) and o3-mini with high reasoning-effort.
Reference-based Reference-free
GPT-4o o3-mini-high GPT-4o o3-mini-high
0.67 0.71 0.67 0.67
Analysis on Human Alignment for Evaluation
We measure their correlation with human judgments for assessing the reliability of model-based evaluation metrics. Table 5 reports the rank correlation between human scores and model-based scores across both reference-based and reference-free settings, using two evaluation models: GPT-4o and o3-mini-high. The results show moderate to strong correlations in all settings. In the reference-based evaluation, o3-mini-high achieves the highest correlation with human judgments (0.67), followed closely by GPT-4o (0.63). In the reference-free setting, both models exhibit strong alignment with human scores, each achieving a correlation of 0.63. These results suggest that model-based evaluations can serve as reliable proxies for human judgment. This supports the use of automatic scoring in large-scale experiments and reinforces the credibility of the evaluation results presented in this study. Based on these findings, we adopt the o3-mini-high model as the evaluation model in all experiments.
@@
To investigate the impact of the underlying language model on overall performance, we experiment with four different LLM backbones: DS-Coder (DeepSeek-Coder-V2-Lite-Instruct), Qwen-Coder (Qwen2.5-Coder-7B-Instruct), DS-Distill-Qwen (DeepSeek-R1-Distill-Qwen-14B), and o3-mini-high (the high reasoning-effort variant of o3-mini). Each model is used within the our framework, and we evaluate the outputs using reference-based, reference-free, and human evaluation metrics. Results are summarized in Table 4. We observe that o3-mini-high consistently outperforms all other models across all evaluation dimensions. In terms of reference-based and reference-free scores, it achieves 3.74 and 4.71, respectively, with a higher score of 4.79 in human evaluation. Notably, it positions the top ranking in all settings, with an average rank of 1.00 in both reference-based and reference-free metrics, and 1.15 in human evaluations. Among the other models, DS-Distill-Qwen performs best, followed by Qwen-Coder and DS-Coder. This performance trend holds consistently across both model-based scores and human judgments. These results highlight the importance of selecting a capable and well-aligned backbone model for generating high-quality, executable repositories from research papers. Based on this results, we mainly use the o3-mini-high as the basis backbone model for all experiments.
Table 6: Ablation study results on PaperCoder. We report the average scores and standard deviations on the ICML 2024 subset of Paper2Code benchmark using model-based reference-based and reference-free evaluation.
@@
Ablation Studies
We conduct ablation experiments to examine the contribution of each module in our multi-stage pipeline. Starting with a base that uses only the full paper text (Paper), we incrementally introduce components in the order they are executed: overall plan, architecture design, core logic, configuration file, and finally the analysis module. As shown in Table 6, performance steadily improves as additional steps are incorporated. Performance temporarily dips when the architecture design module is added. This can be attributed to the lack of dependency-aware ordering at this stage. For example, generating main.py before defining utils.py, which is only resolved later during the core logic phase. Once core logic is introduced, which defines file dependencies and generation order, scores improve substantially. Subsequent additions, such as the configuration file and final analysis stage, further refine the output. Our full system (Ours) achieves the highest scores under both reference-based (3.72) and reference-free (4.73) evaluation settings. These findings highlight the importance of sequential structuring in our pipeline.

### v3

@@ 3.2.3 Coding
The final stage of our PaperCoder framework is the coding phase, where the complete code repository is produced. In particular, each file is generated based on all the available contextual information accumulated from the previous stages, including the overall plan, architecture design, logic design, configuration file, and file-specific analyses, as well as the original paper. Additionally, to ensure consistency across different files, we generate them sequentially according to the execution order (i.e., the ordered file list determined during the logic design stage). To be formal, the overall code generation process is defined as follows: for each file $f_{i}$, the corresponding code $c_{i}$ is generated as $M_{\text{code}}(R,P,f_{i},a_{i},\{c_{1},...,c_{i-1}\}):=\texttt{LLM}(\mathcal{T}_{\text{code}}(R,P,f_{i},a_{i},\{c_{1},...,c_{i-1}\}))\rightarrow c_{i}$, resulting in the complete code repository $C=\{c_{i}\}_{i=1}^{n=|F|}$. We note that this iterative formulation can ensure that $i$-th code is generated with full awareness of its dependencies and the evolving state of the repository.
@@ 4 Experimental Setup
We now describe the experimental setup, including datasets, methods, and evaluation metrics.
4.1 Datasets
To validate the efficacy of our PaperCoder framework in transforming machine learning papers into their corresponding repository-level codes, we construct a new benchmark dataset, which we refer to as Paper2CodeBench. Specifically, we collect the accepted papers from recent machine learning venues (such as ICLR, ICML, and NeurIPS 2024) with the OpenReview API [^1] https://docs.openreview.net/reference/api-v2, and filter them based on the availability of code with its total number of tokens less than 70,000, to ensure the full repository remains within reasonable processing limits of modern LLMs for generation and evaluation. Also, to maintain the benchmark quality, we perform model-based evaluation [22] with GPT-4o on all the collected repositories and select the top 30 from each venue, resulting in a total of 90 papers listed in Tables 1112, and 13. Also, we additionally consider 21 papers for human evaluation (See Table 15). In addition to Paper2CodeBench, we also use the recently released PaperBench Code-Dev [38], which consists of 20 papers from ICML 2024 with paper-specific rubrics annotated by humans. In particular, those rubrics are used to judge the correct implementation based on LLM-based evaluation.
@@ 4.2 Baselines and Our Model
We target the novel problem of Paper2Code, and there are no baselines designed for it to enable direct comparison. Nevertheless, we consider several related approaches proposed to implement repository-level code (or the entire software) from natural language inputs (such as software requirements), in addition to the ablated variants of our full PaperCoder framework, as follows: ChatDev [34] is a multi-agent framework for software development, where several role-specific LLM-powered agents collaborate via structured dialogues; MetaGPT [13] similarity adopts a role-based multi-agent paradigm, but its development process is organized by the principle of Standardized Operating Procedures (SOPs); Abstract is a naive variant of our PaperCoder, which uses only the paper abstract for implementation; Paper, while using the full paper, prompts LLMs only once for code generation; PaperCoder (Ours) is our full framework, structured into three stages of planning, analysis, and code generation. In addition to them, for the PaperBench Code-Dev, we consider baselines suggested by it: Basic Agent is the agentic architecture that can run a predefined set of tools with the ReAct-style approach [49], built upon the basic agent from Inspect AI [^2] https://inspect.ai-safety-institute.org.uk/agents.html#sec-basic-agent, and Iterative Agent that extends Basic Agent, iteratively instructing the model to complete the next subtask.
4.3 Evaluation Setup
Recall that, as shown in Figure 1, the official code implementations of many papers are not available; however, manually annotating their corresponding code implementations to evaluate the quality of automatically generated code repositories is highly labor-intensive and challenging. To address this and ultimately perform the evaluation at scale, we design two model-based evaluation protocols: reference-based (when ground-truth code is available) and reference-free (when it is not), following the recent trends in using LLMs to judge the quality of model outputs [52, 11, 22], especially when they are long. In addition to this, we also perform human evaluations with the authors of the original papers, to ensure reliable judgments and to assess the quality of our model-based evaluations by measuring their correlation with human scores. We discuss each evaluation protocol in detail below.
@@ 4.3 Evaluation Setup
In this evaluation setup, we use the official author-released repository as the gold standard if it is available, since it most accurately reflects the implementations intended by the authors, including the components they consider essential for reproducing the main results presented in the paper. Specifically, we prompt the evaluation model (such as o3-mini-high [^3] Unless otherwise stated, we use o3-mini-high due to strong code understanding and reasoning capability.) to judge the quality of the generated repository with respect to the gold repository, alongside the input paper as context (See Appendix E for the detailed prompt). During evaluation, the model then identifies key components (to be implemented), categorizes them into three severity levels (high, medium, and low), and critiques how well each component is implemented. After that, it returns the overall score on a 5-point Likert scale. We note that, to ensure the reliability of the model-based evaluation, we sample the outputs multiple times (for example, 8 generations per evaluation) and report the averaged score.
@@
For cases where the official author-released code is not available, we introduce the reference-free evaluation protocol that leverages only the paper to assess the quality of its generated repository. In other words, similar to the reference-based evaluation protocol, the evaluation model is prompted to identify key components, categorize them by severity, and critique their implementations in the generated code, but they are performed solely based on the information provided in the paper without access to the ground-truth repository. The rest of the evaluation process, such as sampling and score averaging, follows the same setup as in the reference-based evaluation.
@@
While model-based evaluation offers a scalable and automated way of assessment, we also conduct human evaluations to validate our PaperCoder based on the expert-grounded evaluation. Specifically, to ensure informed and accurate judgment, each participant is assigned a paper for which they are the first author. Furthermore, they are presented with multiple implementations generated by different approaches, and asked to rank them. We provide more details in Appendix A.2.
Evaluation on PaperBench Code-Dev
For evaluation on PaperBench Code-Dev [38], we follow their evaluation setup, measuring the score over the paper-specific rubrics with LLM-based evaluation.
Table 1: Results on our Paper2CodeBench, where we report average scores and standard deviations (in parentheses) grouped by conferences. Oracle denotes the evaluation results with the official repository released by the paper authors. Also, on the right side, we report statistics on the number of tokens, files, and functions, averaged over all implementations. Bold indicates the best scores, statistically significant than baselines ($p\leq 0.05$).
Reference-Based Evaluation Reference-Free Evaluation Statistics
ICLR ICML NeurIPS ICLR ICML NeurIPS # of Tokens # of Files # of Funcs
ChatDEV 2.70 (0.63) 2.97 (0.58) 2.96 (0.69) 4.00 (0.65) 4.12 (0.53) 4.01 (0.74) 6150.54 6.99 23.82
MetaGPT 2.48 (0.48) 2.75 (0.70) 2.95 (0.87) 3.52 (0.60) 3.63 (0.75) 3.59 (0.92) 5405.21 3.24 18.08
Abstract 2.28 (0.42) 2.43 (0.49) 2.35 (0.62) 3.03 (0.64) 3.01 (0.60) 2.99 (0.78) 3376.99 1.28 12.62
Paper 3.08 (0.66) 3.28 (0.67) 3.22 (0.80) 4.15 (0.63) 4.30 (0.53) 4.08 (0.84) 3846.33 1.79 14.84
PaperCoder 3.68 (0.52) 3.72 (0.54) 3.83 (0.50) 4.73 (0.32) 4.73 (0.44) 4.77 (0.38) 14343.38 6.97 35.22
Oracle N/A N/A N/A 4.84 (0.26) 4.80 (0.32) 4.83 (0.38) 32149.04 28.00 122.03
@@ 5 Experimental Results and Analysis
Table 1 presents main results on Paper2CodeBench, from which we observe that PaperCoder consistently outperforms all baselines on both the reference-based and reference-free evaluation settings by substantial margins. We hypothesize that this performance gap of the proposed PaperCoder stems from its top-down behavior, analyzing full papers thoughtfully before starting code generation, unlike prior approaches that typically follow the bottom-up strategy, which begins with and expands short requiremental descriptions (via role-playing or Standardized Operating Procedures). In other words, the top-down approach, operationalized within PaperCoder through the sequential process of planning, analysis, and coding, is effective in handling long-form scientific documents, which are often loosely structured from a software engineering perspective (See Table 6 for more analyses). Also, when compared to the non-comparable Oracle setting (which performs evaluations on the author-released code repositories), PaperCoder achieves performance that is on par, with only marginal differences without a statistically significant gap, demonstrating the effectiveness of our approach in faithfully implementing code whose quality is closer to the implementation by authors.
Correlation between Reference-Based and Reference-Free Evaluation
Recall that the reference-free evaluation protocol is designed for cases where the ground-truth repository is not available, and to investigate whether it works as a reliable proxy for the reference-based evaluation protocol, we measure their rank correlation on all samples from Paper2CodeBench. Then, as shown in Figure 3, there is a strong positive correlation between the reference-based and reference-free evaluations, achieving a Pearson correlation coefficient of $r=0.79$. This result supports that the reference-free evaluation (despite the absence of ground-truth repositories) can serve as a reliable proxy for the reference-based evaluation, ultimately functioning as a standalone metric to assess the code quality.
@@
Figure 3: Correlation between model-based evaluations: reference-based and reference-free.
Table 2: Results with human evaluation. For model-based evaluations (both reference-based and reference-free), 5-point Likert evaluation scores are converted to rankings for comparability with human ranking results. Human rankings are also converted to scores of 5 (top repository), 3 (middle repository), and 1 (bottom repository).
@@
Abstract 2.26 (0.37) 2.94 (0.61) 2.68 (0.56) 2.96 (0.20) 2.96 (0.00) 2.70 (0.56)
Paper 3.00 (0.54) 3.91 (0.63) 2.76 (1.20) 1.92 (0.41) 1.88 (0.38) 2.09 (0.60)
PaperCoder (Ours) 3.66 (0.43) 4.55 (0.51) 4.60 (1.00) 1.08 (0.28) 1.08 (0.28) 1.22 (0.52)
ChatDEV 2.68 (0.60) 3.82 (0.37) 2.12 (1.17) 2.58 (0.50) 2.23 (0.59) 2.43 (0.59)
MetaGPT 2.61 (0.54) 3.39 (0.67) 2.12 (1.17) 2.38 (0.58) 2.46 (0.51) 2.43 (0.59)
PaperCoder (Ours) 3.66 (0.43) 4.55 (0.51) 4.76 (0.88) 1.04 (0.20) 1.04 (0.20) 1.13 (0.46)
Table 3: PaperBench Code-Dev results. We report the averaged performance over three runs with standard deviations.
Replication Score (%)
Model o3-mini-high claude-3.5-sonnet
BasicAgent $5.1\pm 0.8$ $35.4\pm 0.8$
IterativeAgent $16.4\pm 1.4$ $27.5\pm 1.6$
PaperCoder $\textbf{45.14}\pm 0.3$ $\textbf{51.14}\pm 1.4$
Table 4: Results based on both model-based and human evaluations with varying backbone LLMs for PaperCoder.
DS-Coder Qwen-Coder DS-Distill-Qwen o3-mini-high
Score ($\uparrow$) Ref-based 1.47 (0.46) 1.78 (0.28) 2.05 (0.25) 3.66 (0.43)
Ref-free 1.62 (0.54) 2.09 (0.22) 2.31 (0.24) 4.55 (0.51)
Human 1.32 (0.58) 2.71 (1.12) 3.29 (0.98) 4.68 (0.80)
Ranking ($\downarrow$) Ref-based 3.46 (0.00) 2.92 (0.88) 2.25 (0.65) 1.00 (0.20)
Ref-free 3.50 (0.00) 2.88 (0.83) 2.12 (0.54) 1.00 (0.25)
Human 3.74 (0.45) 2.74 (0.86) 2.30 (0.70) 1.22 (0.60)
Table 5: Rank correlation coefficient between human and model-based evaluations (with GPT-4o or o3-mini).
GPT-4o o3-mini-high
Ref-based 0.74 0.78
Ref-free 0.71 0.73
@@
In addition to automatic model-based evaluations, we conduct human evaluations and report the results in Table 2. From this, we confirm that PaperCoder achieves the best ranking among all compared methods, alongside the best scores with model-based evaluations, which reaffirms its effectiveness. Moreover, to ensure whether the model-based evaluations are a reasonable proxy to judge the implementation quality for papers, we measure their correlations with human evaluation scores. As shown in Table 5, we observe strong rank correlations across both reference-based and reference-free settings (where GPT-4o and o3-mini-high are used as evaluation models), which suggests that model-based evaluation can reliably approximate human judgment. Also, based on this result, we use o3-mini-high as the default evaluation model in our experiments due to its highest correlation. Lastly, we ensure the quality and reliability of human evaluations by measuring the inter-annotator agreement score based on Cohen’s kappa coefficient, which exhibits a high score of 0.79, indicating strong consistency.
Results on PaperBench Code-Dev
In addition to our Paper2CodeBench, we further validate the effectiveness of PaperCoder on another PaperBench Code-Dev dataset, which enables fine-grained evaluations for code implementations. As Table 3 shows, PaperCoder achieves the highest replication scores across two different LLMs of o3-mini-high and Claude 3.5 Sonnet, substantially outperforming baselines designed for PaperBench Code-Dev. These results further demonstrate the generalizability and robustness of PaperCoder across diverse evaluation benchmarks and models.
@@
Extending the model variations results on PaperBench Code-Dev, we conduct an auxiliary analysis with DS-Coder (DeepSeek-Coder-V2-Lite-Instruct [8]), Qwen-Coder (Qwen2.5-Coder-7B-Instruct [15]), DS-Distill-Qwen (DeepSeek-R1-Distill-Qwen-14B [9]), and o3-mini-high (the high reasoning-effort variant of o3-mini) on Paper2CodeBench. As summarized in Table 4, the proprietary model (o3-mini-high) consistently outperforms all other backbones across all evaluation settings. Among other open-source models, DS-Distill-Qwen performs the best, followed by Qwen-Coder and DS-Coder. These results suggest the importance of selecting a capable backbone to instantiate PaperCoder, particularly one with strong reasoning capabilities. Also, based on this result, we primarily use o3-mini-high as the basis backbone model for all experiments.
Fine-Grained Analysis of Generated Repositories
To more thoroughly evaluate the quality and practical utility of the generated code, we conduct a set of fine-grained human analyses, according to two perspectives: its usability in reproducing the original work and its component-wise implementation quality. Specifically, we first ask the human annotators whether the top-ranked repository generated by PaperCoder would make it easier to reproduce the methods and experiments compared to starting from scratch. Then, a majority of participants (92%) agree that it would facilitate reproduction, reinforcing the real-world usefulness of our approach. Also, we conduct a component-level implementation analysis to assess which parts of the papers are most effectively translated into code, by asking human annotators to identify the key component-specific elements for Data Processing, Method, and Evaluation, and then measure how many of these elements are actually implemented in the generated code. As shown in Figure 5, we find that the sections of Method, Evaluation, and Data Processing achieve the highest coverage scores, in that order, with each exceeding 50%. Notably, among the errors observed, many of them stem from missing or incorrect implementation details, often rooted in uncertainties about the structure and format of the data, as well as how the data is loaded, which are often under-specified in the papers. Lastly, to investigate why human annotators prefer PaperCoder over its baselines and ablated variants (with 22 out of 25 selecting the repositories from PaperCoder), we ask them to provide the reasons for their choices, and the majority of which are completeness, clean structure, and faithfulness to the original papers, summarized in Table 10.
Ablation Studies
To see how much each component of PaperCoder contributes to the performance gain, we conduct ablation studies on the subset of Paper2CodeBench (composed of ICML papers). Specifically, we start with the base method that uses only the full paper (called Paper) and incrementally add components in the order they are executed (such as overall plan, architecture design, logic design, configuration generation, and final analysis), whose results are reported in Table 6. From this, we observe that the performance steadily improves as additional components are incorporated. Also, we observe a performance drop when the architecture design module is added; however, while this might seem surprising at first, it is in fact expected: architecture design alone does not specify the execution or implementation order of files, which leads to confusion during the code generation stage. However, this issue is effectively addressed once the subsequent logic design module explicitly defines file dependencies and establishes a clear generation order, leading to a substantial performance boost. Overall, integrating all modules in the pipeline yields the highest performance, confirming the effectiveness of our fully structured, multi-stage pipeline, with each module playing a valuable role.
Correlation on Paper Type
To see whether the acceptance category (or presentation format) of papers correlates with the quality of their corresponding implementations by PaperCoder, we analyze it by separating papers into oral/spotlight and poster categories on Paper2CodeBench (which includes 14 oral or spotlight papers and 76 poster papers). As shown in Figure 5, scores are slightly higher for oral/spotlight papers on model-based evaluations with GPT-4o and o3-mini, suggesting that papers with higher recognition might reflect clearer writing, probably leading to faithful code generation.
Figure 4: Detailed, fine-grained analyses on code by PaperCoder.
Table 6: Ablation results on the subset of Paper2CodeBench (ICML papers) with scores and standard deviations.
@@
Figure 5: Model-based evaluation results by presentation type of papers (poster vs. oral/spotlight).